OpenStudy (anonymous):

PLEASE HELP 1.Supplementary angles are two angles whose measures have a sum of 180 degrees. Angles X and Y are supplementary, and the measure of X is 24 degrees greater than the measure of angle Y. Find the measure of each angle. a. X = 102, Y = 78 b. X = 78, Y = 102 c. X = 156, Y = 24 d. X = 24, Y = 156 2. If 2x - 3y = -9 and 3x - 3y = -12, what is the value of y in the solution? a. -3 b. 1 c. (-3, 1) d. (1, -3) 3.What is the solution of 4x + 2y = 8 and 2x + 2y = 2? a. (-2,3) b. (3, 2) c. (3, -2) d. (12, -3)

OpenStudy (anonymous):

please help me @adamarbour

OpenStudy (anonymous):

what you do is 180-24=x then x/2=x+y then to get the two answers it will be (x+24)+y=180 oh and the x and y are just the numbers you should get its not an equation or anything

OpenStudy (anonymous):

thats #1

OpenStudy (mathstudent55):

Problem 1. The problem states: "Supplementary angles are two angles whose measures have a sum of 180." "Angles X and Y are supplementary" From these two statements you can write the equation: X + Y = 180 " the measure of X is 24 degrees greater than the measure of angle Y." From this statement, you can write the equation: X = Y + 24 Here are the two equations together: X + Y = 180 X = Y + 24 With the two equations above, you have a system of equations that can be solved for X and Y. Since the second equation is already solved for X, my suggestion is use the substitution method.

OpenStudy (mathstudent55):

Problem 2. 2x - 3y = -9 3x - 3y = -12 Since you want to solve for y, you need to eliminate x. Multiply the first equation by -3. You will get -6x for the x-term. Multiply the second equation by 2. You will get 6x for the x-term. Add the equations to eliminate x, and solve for y.

OpenStudy (mathstudent55):

4x + 2y = 8 2x + 2y = 2 Subtract the second equation from the first equation. The y-terms will be eliminated, and you can solve for x. Then substitute the value of x into the first equation, and solve for y.

OpenStudy (anonymous):

ok